# Multi-Source EEG Emotion Recognition via Dynamic Contrastive Domain Adaptation

### v1

@@ Abstract
Electroencephalography (EEG) provides reliable indications of human cognition and mental states. Accurate emotion recognition from EEG remains challenging due to signal variations among individuals and across measurement sessions. To address these challenges, we introduce a multi-source dynamic contrastive domain adaptation method (MS-DCDA), which models coarse-grained inter-domain and fine-grained intra-class adaptations through a multi-branch contrastive neural network and contrastive sub-domain discrepancy learning. Our model leverages domain knowledge from each individual source and a complementary source ensemble and uses dynamically weighted learning to achieve an optimal tradeoff between domain transferability and discriminability. The proposed MS-DCDA model was evaluated using the SEED and SEED-IV datasets, achieving respectively the highest mean accuracies of $90.84\%$ and $78.49\%$ in cross-subject experiments as well as $95.82\%$ and $82.25\%$ in cross-session experiments. Our model outperforms several alternative domain adaptation methods in recognition accuracy, inter-class margin, and intra-class compactness. Our study also suggests greater emotional sensitivity in the frontal and parietal brain lobes, providing insights for mental health interventions, personalized medicine, and development of preventive strategies.
Index Terms:
Domain adaptation, unsupervised learning, contrastive learning, dynamic learning, brain-computer interface, EEG feature, emotion recognition, multi-source
I Introduction
@@
Figure 1: Comparisons between conventional discrepancy-based domain adaptation methods with our model. (a) Single-source domain adaptation (SS-DA) treats data from different subjects as a single source (S) and aligns target (T) with S, ignoring the non-stationarity among individual sources. (b) Multi-source domain adapation (MS-DA) align T with each individual source (e.g., S1, S2, S3), tend to produce sub-optimal results due to the lack of fine-grained alignment and trade-offs between domain transferability and discriminability. (c) Our multi-source dynamic contrastive domain adaptation (MS-DCDA) model adapts T to individual sources and a complementary multi source ensemble (SE) with class-awareness. It also dynamically adjusts the weights of domain transferability and discriminability, leading to improved classification accuracies, wider inter-class margin, and higher intra-class compactness.
@@
Unsupervised domain adaptation transfers knowledge from the source domain to the target domain, thereby minimizing the need for extensive data labeling [14, 15, 16]. Unlike single-source domain adaptation (SS-DA), which treats EEG data from different subjects as a single source (see Fig. 1a), multi-source domain adaptation (MS-DA) considers each subject as an individual source domain (see Fig. 1b). The diverse data distribution assumed in the MS-DA may reduce domain bias and model overfitting [17, 18]. However, domain shifts occur not only between each source and target, but also among different sources in the multi-source case, potentially complicating the learning. Two common MS-DA include discrepancy-based and adversarial-based methods [19]. The adversarial discriminative methods align target and source features through domain discriminator and adversarial objectives, whereas the discrepancy-based methods align features in latent space explicitly through discrepancy measures. Existing MS-DA approaches tend to assume shared features across domains, and their domain discriminability is limited to such coarse-grained global alignment. In addition, the typical statitc loss-weighting tends to produce suboptimal balance between the domain alignment and domain discriminability.
@@
To address these challenges, we present a multi-source dynamic contrastive domain adaptation (MS-DCDA) method for EEG-based emotion recognition. Our model learns domain knowledge from both the source and target, where not only we consider the data contribution of each individual source but also a complementary source ensemble (see Fig. 1c). Additionally, our model also learns domain-variant features using fine-grained class-aware contrastive sub-domain adaptation. What is more, an optimal domain transferability and domain discrimination are achieved through a dynamically weighted loss function. Extensive evaluation of the proposed MS-DCDA is conducted on the SEED and SEED-IV dataset, and state-of-the-art performances are achieved for the cross-subject and the cross-session recognition respectively. In addition, the generalization performance of our model is evaluated through a dataset transfer study and key brain lobes involved in EEG emotion recognition are examined. The specific contributions of our work are summarized as follows:
@@
  • Our model leverages domain knowledge from each individual source and a complementary source ensemble uses dynamically weighted learning to achieve optimal domain transferability and discriminability.
  • Our model achieves state-of-the-art performance for cross-subject and cross-session EEG emotion recognition on the SEED and SEED-IV dataset in terms of accuracy, inter-class margin, and intra-class compactness.
II RELATED WORK
II-A EEG-based emotion recognition
@@
EEG provides a non-invasive measurement of the brain’s electrophysiological activities. EEG signals primarily originate from cortical pyramidal neurons in the cerebral cortex, which are oriented perpendicularly to the brain’s surface. These signals offer numerous valuable clinical indications, including those related to higher cognitive functions such as emotions[20]. Two types of commonly used features for emotion classification include the time-domain features (e.g., standard deviation, mean, variance) and the frequency-domain features (e.g., spectral power, differential entropy, and power spectral density). In the frequency domain, the use of differential entropy has demonstrated higher accuracy and greater stability in emotion classification than the power spectral density [21]. Traditional machine learning methods such as support vector machine [22, 23, 24, 25], K-nearest neighbor [26], and random forest [27] make use of these manually crafted features for recognition. In recent years, machine learning and deep learning has achieved remarkable advancement in many areas [28, 29, 30, 31], including affection computing and emotion recognition [10]. Deep learning methods learn features from data without manual feature extractions and selections. For example, convolutional neural networks have been coupled with local information of multiple channels and frequency bands for emotion classification [32]. In addition, an improved emotion classification accuracy of was achived on the SEED dataset by investigating the critical frequency bands and channels via deep belief networks [33].
II-B Single-source domain adaptation
Unsupervised domain adaptation has proven effective in addressing individual differences in EEG signals for emotion recognition. MMD has been explored for cross-session domain adaptation [34]. The deep adaptation network [35] uses MMD to measure the discrepancies between the source and target domain, eliminating individual differences in EEG signals by maintaining discriminative features and domain invariance. Fine-grained information may not be identified using global domain adaptation, and fine-grained adaptation based on sentiment labels achieved better results. A class-aware subdomain adaptation network was introduced [36] based on contrastive domain discrepancy [37], for cross-subject and cross-session emotion recognition. The joint distribution adaptation method [38] approximates the joint distribution by adapting both the marginal and conditional distributions. By making use of emotional labels, improved clustering and recognition accuracy on the SEED dataset is achieved. A deep subdomain adaptation network [39] also capture fine-grained information from each emotion category, improving the alignment of subdomains for domain transfer.
II-C Multi-source domain adaptation
@@
Unified convergent learning bound has been investigated and extended to multi-source data [40]. The deep multi-source adaptation transfer network [41] combines the deep adaptation network [35] and a discriminator, allowing nonuniform distribution in cross-subject emogion recognition. The multi-source marginal distribution adaptation method [21] maps domains to a shared feature space using multi-layer perceptrons to extract unique domain-invariant features in each individual domain. The multi kernel and multi-source MMD method [42] extends the traditional single-source MMD to measure the differences between domains. Joint distribution was investigated in multi-source domain adaptation [43] to establish a fine-grained domain adaptation through conditional distribution of each domain using reinforcement learning, where pseudo labels were employed in class-level domain alignment [44].
@@
Figure 2: The pipeline of EEG-based emotion recognition and the schematics of the proposed MS-DCDA model. The multi-source EEG data are measured and preprocessed. The five-band differential entropy (DE) features are extracted from the data of each participant. The multi-source dynamic contrastive domain adaptation (MS-DCDA) model consists of three modules: the common feature extractor (CFE), the multi-branch contrastive (MBC) module, and the multi-branch domain classifier (MBDC). The CFE module extracts features extract domain-invariant features from source data $\{x_{s}^{i}|i=0,...N\}$ and the target data $x_{t}$ using a shared MLP. The MBC and MBDC modules are respectively comprised of $N+1$ independent branches $\{MBC_{i}\}$ and $\{MBDC_{i}\}$. The MBC module extracts the domain-variant features and the DC module classify them into domain-specific labels. During training, the $\mathcal{L}_{MMD}$ and $\mathcal{L}_{SCD}$ losses encourage class-independent and class-aware alignment respectively. The classification loss is comprised of the cross entropy loss $\mathcal{L}_{CE}$ and a complementary $\mathcal{L}_{DISC}$ loss, which encourages the predictions consistency across classifiers. Additionally, a dynamic coefficient $\tau$ optimizes the domain transferability and discriminability. During the test, the predicted target class probabilities are averaged across the classifiers, and the target emotion $\hat{e}_{t}$ is determined by the maximum mean class probability.
III METHODS
@@
We introduce a multi-source dynamic contrastive domain adaptation (MS-DCDA) method for cross-subject and cross-session EEG-based emotion recognition. Our framework makes use of multi-branch contrastive neural net for class-aware domain alignment, which allows discriminative adjustment of marginal and conditional distributions of domains. The multi-source MMD is combined with a class-aware subdomain contrastive discrepancy (SCD) metric to tradeoff the coarse- and fine-grained domain alignments between source and target domains. Additionally, domain transferability and discriminability are optimally balanced through dynamically weighted domain alignment and classification penalties. The pipeline of EEG-based emotion recognition and the schematic of our MS-DCDA model are illustrated in Fig. 1.
III-A Data Preprocessing
@@
The pipeline of the emotion recognition starts with EEG data preprocessing, where the measured EEG signals are segmented and prefiltered. Differential entropy (DE) [45] features, which are effective in distinguishing low- and high-frequencies in EEG signals, are extracted from the EEG segments of each participant for five frequency bands (see Section IV-A). For an EEG segment having a Gaussian distributed frequency spectrum $\mathcal{N}(\mu,\sigma^{2})$, its DE is given by:
@@ III-B MS-DCDA Architecture
Common Feature Extractor. The common feature extractor (CFE) module extracts domain-invariant features from the source and target domain using a shared neural network. Here we use the precomputed DE features of EEG signals as pre-processed input. The target domain input is denoted as $X_{t}=\{x_{t}\}$, where $x_{t}\in\mathbb{R}^{W\times B}$, $W$ is number of EEG samples per individual subject, $B$ represents the total number of channels, which is the superposition of EEG channels in each frequency band. The source domain consists of $N+1$ sources $X_{s}=\{x_{s}^{i}|i=0,...N\}$. The first $N$ source elements represent the individual sources $x_{s}^{i}\in\mathbb{R}^{W\times B}$ ($i<N$). The last source element indicates an ensemble of the individual sources $x_{s}^{N}=\bigcup_{i=0}^{N-1}\{x_{s}^{i}\}$. $x_{s}^{N}\in\mathbb{R}^{W\times C}$ where $C$ is the total number of channels in the combined domain, which is the total number of channels in the first n-1 source domains. The common target feature is given by $F_{t}=\{f_{t}|f_{t}=CFE(x_{t})\}$. The common source features are extracted not only from each individual source but also from the source ensemble $F_{s}=\{f_{s,D}^{i}|f_{s,D}^{i}=CFE(x_{s}^{i}),i=0,...,N\}$. Here we employ multi-layer perceptron (MLP) for the CFE module, where the structure of each of the three layers is given respectively by Linear310/256/128-BatchNorm1D-LeakyReLU(0.01).
@@
Multi-Branch Domain Classifier. The multi-branch domain classifier (MBDC) module predicts domain-specific labels from the features extracted by the MBC module. The MBDC module is comprised of $N+1$ linear classifiers $\{MBDC_{i}\}$ with structure Linear32-softmax. The predicted labels for each individual source and the ensemble source domain are denoted as $\widehat{Y}_{s}=\{\widehat{y}_{s}^{i}|\widehat{y}_{s}^{i}=MBDC_{i}(f_{s,D}^{i}),i=0,...N\}$. The predicted label of the target is represented as $\widehat{Y}_{t}=\{\widehat{y}_{t}^{i}|\widehat{y}_{t}^{i}=MBDC_{i}(f_{t,D}^{i}),i=0,...N\}$. The ground truth label is given by $Y_{s}=\{y_{s}^{i}|i=0,...,N\}$. During the training, $\hat{Y}_{s}$, $\hat{Y}_{t}$, and $Y_{s}$ are used in learning the domain classfier and serving as class guidances in learning fine-grained domain adaptation. At the prediction (or testing) stage, the emotion state of the target subject is identified from a set of $M$ defined emotions $\{e_{m}|m=1,...,M\}$, each representing an emotion state (e.g., positive, negative, or neutral). The class probabilities of the target domain class $\{p_{t}^{i}=P(\hat{y_{i}}|y_{i})|i=0,...,N\}$ are averaged across the domain classifiers [17], producing the mean class probabilities $\bar{p}_{t}=(\Sigma_{i=0}^{N}p_{t}^{i})/N$, where $\bar{p}_{t}\in\mathbb{R}^{1\times M}$. The emotion state of the target participant is identified by the maximum mean class probability $\hat{e}_{t}=e_{j}$ where $j=\arg\max\limits_{m}\{\bar{p}_{t,m}\}$.
III-C MS-DCDA Learning
The learning of the proposed MS-DCDA model involves four aspects: 1) coarse-grained domain aligment; 2) fine-grained sub-domain domain alignment with class-awareness; and 3) fine-grained classification. 4) Dynamically weighted loss function for optimally balanced domain ility and discrimination performance.
Domain Alignment. The coarse-grained alignment of the source and target involves using the multi-source MMD to measure the difference between the MBC features of the pair-wise source and the target $\{f_{s}^{i},f_{t}^{i}\}$. Mathematically, the MMD loss measures the difference between two distributions using their mean embeddings in the reproducing kernel Hilbert space. In practice, the squared value of MMD is typically estimated by kernel embedding:

### v2

@@ Abstract
Electroencephalography (EEG) provides reliable indications of human cognition and mental states. Accurate emotion recognition from EEG remains challenging due to signal variations among individuals and across measurement sessions. We introduce a multi-source dynamic contrastive domain adaptation method (MS-DCDA) based on differential entropy (DE) features, in which coarse-grained inter-domain and fine-grained intra-class adaptations are modeled through a multi-branch contrastive neural network and contrastive sub-domain discrepancy learning. Leveraging domain knowledge from each individual source and a complementary source ensemble, our model uses dynamically weighted learning to achieve an optimal tradeoff between domain transferability and discriminability. The proposed MS-DCDA model was evaluated using the SEED and SEED-IV datasets, achieving respectively the highest mean accuracies of $90.84\%$ and $78.49\%$ in cross-subject experiments as well as $95.82\%$ and $82.25\%$ in cross-session experiments. Our model outperforms several alternative domain adaptation methods in recognition accuracy, inter-class margin, and intra-class compactness. Our study also suggests greater emotional sensitivity in the frontal and parietal brain lobes, providing insights for mental health interventions, personalized medicine, and preventive strategies.
keywords:
[…]
1 Introduction
@@
Figure 1: Comparison of traditional discrepancy-based domain adaptation methods with our model. (a) Single-source domain adaptation (SS-DA) treats data from different subjects as a single source (S) and aligns target (T) with S, ignoring non-stationarity among individual sources. (b) Multi-source domain adaptation (MS-DA) aligns T with each individual source, but tends to produce sub-optimal results due to the lack of fine-grained alignment and imbalanced domain transferability and discriminability. (c) Our multi-source dynamic contrastive domain adaptation (MS-DCDA) model adapts T to individual sources and a complementary multi-source ensemble (SE) with class-awareness. Dynamically weighted domain transferability and discriminability provides improved accuracies, wider inter-class margin, and higher intra-class compactness.
@@
Domain adaptation transfers knowledge from the source domain to the target domain, thereby minimizing the need for extensive data labeling [14, 15, 16]. Unlike single-source domain adaptation (SS-DA), which treats EEG data from different subjects as a single source (see Fig. 1(a)), multi-source domain adaptation (MS-DA) considers each subject as an individual source domain (see Fig. 1(b)). The assumption of diverse data in the MS-DA may reduce domain bias and model overfitting [17, 18]. However, domain shifts occur not only between each source and target, but also among different sources, potentially complicating the learning. Two common MS-DA include discrepancy-based and adversarial-based methods [19]. The adversarial discriminative methods align target and source features through domain discriminator and adversarial objectives, whereas the discrepancy-based methods align features in latent space explicitly through discrepancy measures. Existing MS-DA approaches tend to assume shared features across domains, and their domain discriminability is limited to such coarse-grained global alignment. In addition, the typical static loss-weighting tends to produce suboptimal balance between the domain alignment and domain discriminability.
@@
To address these challenges, we present a multi-source dynamic contrastive domain adaptation (MS-DCDA) method for EEG-based emotion recognition. Our model learns domain knowledge from both the source and target, where we consider not only the data contribution of each individual source but also a complementary source ensemble (see Fig. 1(c)). Additionally, our model also learns domain-variant features using fine-grained class-aware contrastive sub-domain adaptation. What is more, an optimal domain transferability and domain discrimination are achieved through a dynamically weighted loss function. Extensive evaluation of the proposed MS-DCDA is conducted on the SEED and SEED-IV dataset, and state-of-the-art performances are achieved for the cross-subject and the cross-session recognition respectively. In addition, the generalization performance of our model is evaluated through a dataset transfer study where key brain lobes involved in EEG emotion recognition are also examined. The specific contributions of our work are summarized as follows:
@@
  • Our model leverages domain knowledge from each individual source and a complementary source ensemble uses dynamically weighted learning to achieve optimal domain transferability and discriminability, significant improvements in inter-class margin and intra-class compactness are demonstrated.
  • State-of-the-art performance is achieved on SEED and SEED-IV datasets, and our model outperforms the second best approach respectively by $1.5\%$ and $2.4\%$ in the cross-subject experiments as well as $2.2\%$ and $3.6\%$ in the cross-session experiment.
2 RELATED WORK
2.1 EEG-based emotion recognition
@@
EEG provides a non-invasive measurement of the brain’s electrophysiological activities. EEG signals mainly come from cortical pyramidal neurons perpendicular to the surface of the brain in the cerebral cortex. These signals offer numerous valuable clinical indications, including those related to higher cognitive functions such as emotions [20]. Two types of commonly used features for emotion classification include the time-domain features (e.g., standard deviation, mean, variance, and differential entropy) and the frequency-domain features (e.g., spectral power and power spectral density). In both feature domains, the use of differential entropy has demonstrated higher accuracy and greater stability in emotion classification than the power spectral density [21]. Traditional machine learning methods such as support vector machine [22, 23, 24, 25], K-nearest neighbor [26], and random forest [27] make use of these manually crafted features for recognition. In recent years, machine learning and deep learning has achieved remarkable advancement in many areas [28, 29, 30, 31, 32], including affection computing and emotion recognition [10]. Deep learning methods learn features from data without manual feature extractions and selections. For example, convolutional neural networks [33] have proven effective for lesion classification. CNNs have also been coupled with local information for emotion classification. By exploring frequency bands and channels via deep belief networks [34], the accuracy has been improved on the SEED dataset. Additionally, using attention-based convolutional recurrent neural networks [35] for discriminative features extraction further boosting the performance of EEG based emotion recognition. Pre-trained Transformers are also optimized through random feature projection and head augmentation [36], demonstrating improved efficiency, accuracy, and generalization in classification tasks.
2.2 Single-source domain adaptation
Domain adaptation is a type of transductive transfer learning [16], which adapts a model learned on source domain to classify data in target domain. Depending on the availability of class labels of samples in target domain during the training, deep visual domain adaptation are typically categorized into: supervised, semi-supervised, and unsupervised approaches. The case where the label is unavailable in target domain is considered as unsupervised domain adaptation (UDA) [37] [38][39]. Unsupervised domain adaptation has proven effective in addressing individual differences in EEG signals for emotion recognition, even in cases where the distribution of source and target domains are entirely different. To transfer cross-subject knowledge, maximum mean discrepancy (MMD) [40] has been widely adopted as a metric to describe the discrepancies between the distributions of source and target domains in UDA. For example, MMD preserves the domain invariance in deep adaptation network [41] while addressing the individual differences. However, the naive MMD is not capable of addressing the class prior distributions, resulting in degraded domain adaptation accuracies. Class-aware subdomain adaptation network have been introduced [42, 43] to address the class weight bias. In addition, contrastive domain discrepancy metric [44] is introduced for cross-subject and cross-session emotion recognition. The approximated joint marginal and conditional distributions [45] also shows improved clustering and recognition accuracy on the SEED dataset. A deep subdomain adaptation network [46] captures fine-grained information from each class, improving the alignment of subdomains.
2.3 Multi-source domain adaptation
@@
Single-source DA methods are limited to single-source domain distributions and relatively simple datasets. Multi- source domain adaptation algorithms are introduced to address more complicated data distributions. Unified convergent learning bound has been investigated for multi-source data [47]. The deep multi-source adaptation transfer network [48] combines an adaptation network [41] with a discriminator, allowing nonuniform distribution in cross-subject emotion recognition. The multi-source marginal distribution adaptation method [21] maps domains to a shared feature space using multi-layer perceptrons to extract unique domain-invariant features from each individual domain. The multi-kernel and multi-source MMD method [49] extends the single-source MMD to measure the differences between domains. Joint distribution was also investigated in multi-source domain adaptation [50] for fine-grained domain adaptation using reinforcement learning. Dynamic transfer learning [51] is also explored to reduce multi-source to single-source adaptation thorough dynamic routing.
@@
Figure 2: The pipeline of EEG-based emotion recognition and the schematics of the proposed MS-DCDA model. The multi-source EEG data are measured and preprocessed. The five-band differential entropy (DE) features are extracted from the data of each participant. The multi-source dynamic contrastive domain adaptation (MS-DCDA) model consists of three modules: the common feature extractor (CFE), the multi-branch contrastive (MBC) module, and the multi-branch domain classifier (MBDC). The CFE module extracts features extract domain-invariant features from source data $\{x_{s}^{i}|i=0,...N\}$ and the target data $x_{t}$ using a shared MLP. The MBC and MBDC modules are respectively comprised of $N+1$ independent branches $\{MBC_{i}\}$ and $\{MBDC_{i}\}$. The MBC module extracts the domain-variant features and the DC module classify them into domain-specific labels. During training, the $\mathcal{L}_{MMD}$ and $\mathcal{L}_{SCD}$ losses encourage class-independent and class-aware alignment respectively. The classification loss is comprised of the cross entropy loss $\mathcal{L}_{CE}$ and a complementary $\mathcal{L}_{DISC}$ loss, which encourages the predictions consistency across classifiers. Additionally, a dynamic coefficient $\tau$ optimizes the domain transferability and discriminability. During the test, the predicted target class probabilities are averaged across the classifiers, and the target emotion $\hat{e}_{t}$ is determined by the maximum mean class probability. We train our model in an unsupervised setting, where the source domain is labeled whereas the labels are unavailable for samples in the target domain.
3 METHODS
@@
We introduce a multi-source dynamic contrastive domain adaptation (MS-DCDA) method for cross-subject and cross-session EEG-based emotion recognition. Our framework makes use of multi-branch contrastive neural net for class-aware domain alignment, which allows discriminative adjustment of marginal and conditional distributions of domains. The multi-source MMD is combined with a class-aware subdomain contrastive discrepancy (SCD) metric to tradeoff the coarse- and fine-grained domain alignments between source and target domains. Additionally, domain transferability and discriminability are optimally balanced through dynamically weighted domain alignment and classification penalties. The pipeline of EEG-based emotion recognition and the schematic of our MS-DCDA model are illustrated in Fig. 2. We learn transferable EEG features in an unsupervised domain adaptation (UDA) setting, where the samples are labeled in the source domain but remain unlabeled in the target domain during the training.
3.1 Data Preprocessing
@@
The pipeline of the emotion recognition starts with EEG data preprocessing, where the measured EEG signals are segmented and prefiltered. Differential entropy (DE) [52] features, which are effective in distinguishing low- and high-frequencies in EEG signals, are extracted from the EEG segments of each participant for five frequency bands (see Section 4.1). For an EEG segment having a Gaussian distributed frequency spectrum $\mathcal{N}(\mu,\sigma^{2})$, its DE is given by:
@@ 3.2 MS-DCDA Architecture
Common Feature Extractor. The common feature extractor (CFE) module extracts domain-invariant features from the source and target domain using a shared neural network. Here we use the precomputed DE features of EEG signals as pre-processed input. The target domain input is denoted as $X_{t}=\{x_{t}\}$, where $x_{t}\in\mathbb{R}^{W\times B}$, $W$ is number of EEG samples per individual subject, $B$ represents the total number of channels, which is the superposition of EEG channels in each frequency band. The source domain consists of $N+1$ sources $X_{s}=\{x_{s}^{i}|i=0,...N\}$. The first $N$ source elements represent the individual sources $x_{s}^{i}\in\mathbb{R}^{W\times B}$ ($i<N$). The last source element indicates an ensemble of the individual sources $x_{s}^{N}=\bigcup_{i=0}^{N-1}\{x_{s}^{i}\}$. $x_{s}^{N}\in\mathbb{R}^{W_{2}\times B}$ where $W_{2}$ is the total number of samples in the combined domain, which is the total number of samples in the first n-1 source domains. The common target feature is given by $F_{t}=\{f_{t}|f_{t}=CFE(x_{t})\}$. The common source features are extracted not only from each individual source but also from the source ensemble $F_{s}=\{f_{s}^{i}|f_{s}^{i}=CFE(x_{s}^{i}),i=0,\ldots,N\}$. Here we employ multi-layer perceptron (MLP) for the CFE module, where the structure of each of the three layers is given respectively by Linear310/256/128-BatchNorm1D-LeakyReLU(0.01).
@@
Multi-Branch Domain Classifier. The multi-branch domain classifier (MBDC) module predicts domain-specific labels from the features extracted by the MBC module. The MBDC module is comprised of $N+1$ linear classifiers $\{MBDC_{i}\}$ with structure Linear32-softmax. The predicted labels for each individual source and the ensemble source domain are denoted as $\widehat{Y}_{s}=\{\widehat{y}_{s}^{i}|\widehat{y}_{s}^{i}=MBDC_{i}(f_{s,D}^{i}),i=0,...N\}$. The predicted label of the target is represented as $\widehat{Y}_{t}=\{\widehat{y}_{t}^{i}|\widehat{y}_{t}^{i}=MBDC_{i}(f_{t,D}^{i}),i=0,...N\}$. The ground truth label is given by $Y_{s}=\{y_{s}^{i}|i=0,...,N\}$. During the training, $\hat{Y}_{s}$, $\hat{Y}_{t}$, and $Y_{s}$ are used in learning the domain classifier and serving as class guidances in learning fine-grained domain adaptation. At the prediction (or testing) stage, the emotion state of the target subject is identified from a set of $M$ defined emotions $\{e_{m}|m=1,...,M\}$, each representing an emotion state (e.g., positive, negative, or neutral). The class probabilities of the target domain class $\{p_{t}^{i}=P(\hat{y_{i}}|y_{i})|i=0,...,N\}$ are averaged across the domain classifiers [17], producing the mean class probabilities $\bar{p}_{t}=(\Sigma_{i=0}^{N}p_{t}^{i})/N$, where $\bar{p}_{t}\in\mathbb{R}^{1\times M}$. The emotion state of the target participant is identified by the maximum mean class probability $\hat{e}_{t}=e_{j}$ where $j=\arg\max\limits_{m}\{\bar{p}_{t,m}\}$.
3.3 MS-DCDA Learning
The learning of the proposed MS-DCDA model involves four aspects: 1) coarse-grained domain aligment; 2) fine-grained sub-domain domain alignment with class awareness; and 3) fine-grained classification. 4) Dynamically weighted loss function for optimally balanced domain transferability and discrimination performance.
Domain Alignment. The coarse-grained alignment of the source and target involves using the multi-source MMD to measure the difference between the MBC features of the pair-wise source and the target $\{f_{s,D}^{i},f_{t,D}^{i}\}$. Mathematically, the MMD loss measures the difference between two distributions using their mean embeddings in the reproducing kernel Hilbert space. In practice, the squared value of MMD is typically estimated by kernel embedding: